OpenStudy (annie96):

Question 2.2. Solve the given system, using the substitution method. y = 4x – 6 8x – 2y = 14 (Points : 1) (14, 12) (12, 14) There are an infinite number of solutions. There is no solution. Question 3.3. Solve, using the substitution method. 2x + 3y = 1 y = x – 8 (Points : 1) (0, –8) (2, –6) (4, –4) (5, –3)

jigglypuff314 (jigglypuff314):

y = 4x – 6 <-- plug in this one into 8x – 2y = 14 <-- this one => 8x - 2(4x - 6) = 14 => 8x - 8x + 12 = 14 12 = 14 <-- no solution

jigglypuff314 (jigglypuff314):

2x + 3y = 1 y = x – 8 <-- plug in this one, into above one so 2x + 3(x-8) = 1 distribute 2x + 3x - 24 = 1 add like terms 5x - 24 = 1 add 24 to both sides +24 +24 5x = 25 divide both sides by 5 /5 /5 x = 5 <-- which option had that number in it?